f(x,y)=4x-3y function on the curve x^2+y^2=25 when finding the extreme values using Lagrange multipliers which of these statements are true?

1)f is at minimum value at (-4,-3) on this curve
2)f is at maximum value at (4,3) on this curve
3)maximum value of f is 25 on this curve
4)f has not got any extremum points on this curve
Answer #1

Given, function finger 4x-3y google tye - 25 subjected to solution: Take all will the partial derivatives which for the Lagrawl A 11 ४. ਹੈ - wo eq14) y → 3 4 x into eqh (3) to find: substituting x +39=25 ? 2 น + 9x² = 25 介 16 25ne = 25 16 = > u= 4 N• f14,-3) = 4(4) – 36-3) = 16+9 = 25. • f(-4,3) 41-4) - 3/3) = -16-9= -25. at (4,-3) so, the value = 25 Maximum value = -25 a